OpenStudy (anonymous):

A pole that is 3.3m tall casts a shadow that is 1.48m long. At the same time, a nearby tower casts a shadow that is 49.75m long. How tall is the tower? Round your answer to the nearest meter

OpenStudy (anonymous):

We'll have to make a proportion. Do you know how, or do you need help with that too.

OpenStudy (anonymous):

yeah, i think i can do that. would it me 3.3/1.48 = x/49.75?

OpenStudy (anonymous):

Yes, that looks right to me. Think you can solve it?

OpenStudy (anonymous):

yes, is that the only thing that i need to do, or is there more too it?

OpenStudy (anonymous):

\[\frac{ 3.3 }{ 1.48 }=\frac{ x }{ 49.75 }\]

OpenStudy (anonymous):

That should give you the answer. I put it in that form in case it helps.

OpenStudy (anonymous):

Thank you! i have found the answer. i got 110.929, when i round it to the nearest meter it will be 111m?

OpenStudy (anonymous):

Yes, good job :)

OpenStudy (anonymous):

Thank you so much! i really appreciate it.

OpenStudy (anonymous):

No problem. Good to see someone try by themselves first. :)

OpenStudy (anonymous):

is it okay if i tag you in other questions as well, because i am most likely going to need help.

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

okay, thank you! :)